In 2016, Manhoff Company had a break-even point of $341,000 based on a selling price of $5 per unit and fixed costs of $171,000. In 2017, the selling price and the variable costs per unit did not change, but the break-even point increased to $437,000.

Compute the variable costs per unit and the contribution margin ratio for 2016.

AND

Compute the increase in fixed costs for 2017.
Answer #1

1. Break even sales = {Fixed expense /(sales - variable cost)}* sales

341000 = {171000/(5-variable cost per unit)}*5

Variable cost per unit = 2.51

Contribution margin ratio = ((5-2.51)/5 = 49.8%

2. Increase in break even = 437000-341000 = $96000

Increase in fixed cost = 96000*49.8% = 47808